A 2270 kg safety test vehicle is traveling at 34.4 m/s when it slams into a brick wall. It takes 0.190 seconds for the vehicle to stop. What is the average force exerted on the test vehicle?



1
Expert's answer
2020-12-28T08:41:50-0500

"F=\frac{m\Delta v}{\Delta t}=\frac{2270\cdot 34.4}{0.19}=411(kN)"